An object was dropped from a plane that is moving in the sky horizontally with a speed of 272 m/s. If the object took 26 seconds before it hits the ground, find a) the height from which it was dropped, and b) the speed of the object before it hits the ground.
1
Expert's answer
2018-09-24T12:09:09-0400
a)
h=(gt^2)/2=9.8/2 (26)^2=3312 m.
b)
v_y=gt.
v=√(v_x^2+v_y^2 )=√(〖272〗^2+(9.8∙26)^2 )=373 m/s.